OpenStudy (anonymous):

@hoblos : Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2)

OpenStudy (hoblos):

slope = (-2-8)/(0-4) = 10/4 = 5/2 use the formula y - y1 = m(x-x1) then y-8 = (5/2)(x-4) y - 8 = (5/2)x - 10 y = (5/2)x -2